Can someone please help with the reactions and equation based on this experiment. I am trying to write a lab report

Experiment 14 ates of Reaction: a Clock Reaction Objectives .To determine the overall order of a reaction To determine the effects of concentration, temperature and catalyst on a reaction rate Prior Reading Chemistry The Central Science Related topics in Chapter 14 Introduction Kinetics represents the study of the rates of chemical reactions. The rate at which reactants are converted into products is dependent upon many factors such as the nature of the reaction, the concentration of reactants, the temperature, and the presence of a catalyst. For any given reaction, such as the generic reaction: Experiment 14-1
Can someone please help with the reactions and equ
Can someone please help with the reactions and equ
Can someone please help with the reactions and equ
Can someone please help with the reactions and equ
Can someone please help with the reactions and equ
Can someone please help with the reactions and equ
Answer #1

for the given reaction

rate is given by

Rate = K\left [ I^{-} \right ]^{m}\left [ BrO_{3} ^{-}\right ]^{n}\left [ H^{+} \right ]^{p}

and order of reaction = m+n+p

concentrations of the reactants are calculated by

M_{1}V_{1}=M_{2}V_{2}

the rate law coefficients m,n,p are caculated by keeping any two reactant concentrations constant and varying only one.there by identifying m n and p individually

and finally determining the rate of reaction.

once rate of reaction is obtained activation energy can be calculated by the equation:

K=Ae^{\frac{-E_{a}}{RT}}

factors effecting the rate of reaction

a) when concentrations are altered if reaction rate is doubled then it is first order and if it is quadrapled it is second order

b) temperature effects the reaction rate directly

c) catalyst speedens the reaction rate.